Thinfilms morning report 2/18/05

james haydon jhaydon at snf.stanford.edu
Fri Feb 18 09:17:05 PST 2005


Thinfilms Morning Report 2/18/2005

 

 

Gryphon:            PPC fixed ok to use

Innotec;            ok

STS Dep:            ok

Implanter:            having vacuum and high voltage problems should be 
able to get to it late next week

Metalica:            problem with target tester will try to fix it next week

SCT:                ok

CPD:                ok                   

WbGen-CTB:            ok

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            ok

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok

 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/thinfilms/attachments/20050218/50cbff1f/attachment.html>


More information about the thinfilms mailing list